2x + 5y = -1
x + 2y = 0
can u please show me how to solve this problem step by step

Found 3 solutions by jim_thompson5910, It is costly, addingup:
Answer by jim_thompson5910(35256)

Step 1) Solve for x in equation 2

x+2y = 0
x+2y-2y = 0-2y ... subtract 2y from both sides
x = -2y

-----------------------------------------------------

Step 2) Replace x with -2y in equation 1. Then solve for y

2x + 5y = -1
2(-2y) + 5y = -1 ... replace x with -2y
-4y + 5y = -1
1y = -1
y = -1

-----------------------------------------------------

Step 3) Use the value of y found in step 2 to find the value of x

x = -2y
x = -2*(-1) ... plug in y = -1
x = 2

======================================================================

Final Answer: x = 2 and y = -1. The two values pair up to form the ordered pair or point (x,y) = (2,-1)

Side Note: If you graph the two original equations on the same xy plane, then you'll see two lines crossing at the point (2,-1).

Answer by It is costly(175)
Answer by addingup(3677)

2x + 5y = -1
x + 2y = 0 On this equation, multiply both sides times 2:
-----------------------------------------------------------
2x+4y= 0 Now subtract this from the first equation
------------------------------------------------------
2x + 5y = -1
-
2x + 4y = 0
----------
... y = -1 Now that we know y we can find the value of x. For this we can use either one of the equations, let's use the second:
-----------------------------------------------------------
x + 2y = 0
x + 2(-1)= 0
x + -2= 0
x= 2
----------------------------------------------------------------
Check:
2(2) + 5(-1) = -1
4 + -5 = -1
-1 = -1 This equation checks out
----------------------------------------------------------------
x+2y = 0
2 + 2(-1)
2 + -2 = 0
0 = 0 This equation also checks out. We have the correct answers
J